A sample ofย ๐‘›=18 observations is drawn from a normal population withย ๐œ‡=960 andย ๐œŽ=220. Find each of the following:

A.ย ๐‘ƒ(๐‘‹ >1037)

Probability =ย 

B.ย ๐‘ƒ(๐‘‹ <877)

Probability =ย 

C.ย ๐‘ƒ(๐‘‹>934)

Probability =ย 


1
Expert's answer
2021-10-11T15:09:19-0400

A.

"z=\frac{x-\mu}{\sigma/\sqrt{n}}=\frac{1037-960}{220/\sqrt{18}}=0.08"


"P(X>1037)=1-P(z<0.08)=1-0.53188=0.46812"


B.

"z=\frac{877-960}{220/\sqrt{18}}=-0.09"


"P(X<877)=P(z<-0.09)=0.46414"


C.

"z=\frac{934-960}{220/\sqrt{18}}=-0.03"


"P(X>934)=1-P(z<-0.03)=1-0.48803=0.51197"